An offshore oil well is located in the ocean at a point 𝑾, which is 5 π‘šπ‘–π‘™π‘’π‘  from the closest shore-point 𝑨 on a straight shoreline. The oil is to be piped to a shore-point 𝑩 that is 8 π‘šπ‘–π‘™π‘’π‘  from 𝑨 by piping it on a straight line underwater from 𝑾 to some shore-point 𝑷 between 𝑨 and 𝑩 and then on to 𝑩 via a pipe along the shoreline. If the cost of laying pipe is $100,000.00 per mile under water and $75,000.00 per mile over land, where should point 𝑷 be located to minimize the cost of laying the pipe?


1
Expert's answer
2021-09-30T23:37:31-0400



Cost of pipe = Cost under water + Cost on land

"=100000x+75000y"


Using Pythagoras theorem, we get:

"(5)^2+(8-y)^2=x^2"

We need to change this equation in terms of one variable:

"=100000(\sqrt{(5)^2+(8-y)^2})+75000y\\"

We need to find the cost price and then minimize it.

Let "C(y)=100000(\sqrt{(5)^2+(8-y)^2})+75000y"

"\therefore C'(y)=100000\times \frac{1}{(\sqrt{(5)^2+(8-y)^2})}\times (-2(8-y))+75000"

Put "C'(y)=0" , we get:

"\Rightarrow100000\times \frac{1}{(\sqrt{(5)^2+(8-y)^2})}\times (-2(8-y))+75000=0"

"\Rightarrow \frac{-200(8-y)}{(\sqrt{(5)^2+(8-y)^2})}+75=0\\
\Rightarrow -200(8-y)=-75{(\sqrt{(5)^2+(8-y)^2})}\\
\Rightarrow 8(8-y)=3{(\sqrt{(5)^2+(8-y)^2})}\\"

Squaring both sides, we get:

"64(8-y)^2=9(25+(8-y)^2)\\
\Rightarrow 55(8-y)^2=225\\
\Rightarrow (8-y)^2=\frac{45}{11}\\
\Rightarrow y=5.97"

So, "8-y=8-5.97=2.03" miles right of point A.